OpenStudy (anonymous):

Find an equation for the circle centered at (1, –5) and tangent to the line 3x + 4y = 8.

OpenStudy (sirm3d):

you can use the formula for the distance of the point (center) from a line (3x+48=8). this distance is teh radius of the circle. you do know the formula?

OpenStudy (anonymous):

can u give me a drawing of it like how it looks

OpenStudy (aripotta):

(x - 1)^2 + (y + 5)^2 = 25

OpenStudy (aripotta):

oh wait!

OpenStudy (aripotta):

(x - 1)^2 + (y + 5)^2 = 25.16

OpenStudy (anonymous):

i know the answer i just need the drawing

OpenStudy (anonymous):

THANK U :)

OpenStudy (aripotta):

no problem :)

OpenStudy (sirm3d):



OpenStudy (anonymous):

thank u @sirm3d

OpenStudy (sirm3d):

YW